40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 19 19:08:40 2011, Jenny, Update, PSL, Relocking NPRO to reference cavity.  
    Reply  Fri Aug 19 23:41:20 2011, Jenny, Update, PSL, Relocking NPRO to reference cavity.  
Message ID: 5272     Entry time: Fri Aug 19 23:41:20 2011     In reply to: 5271
Author: Jenny 
Type: Update 
Category: PSL 
Subject: Relocking NPRO to reference cavity.  

Quote:

I am trying again to measure a temperature step response on the reference cavity on the PSL table.

I have been working to relock the NPRO to the cavity. I unblocked the laser beam, reassembled the setup described in my previous elog entry: 5202. I then did the following:

1) Monitored error signal (from LB1005 PDH servo), transmitted signal, and control signal sent to drive PZT on oscilloscope.

2) With loop open, swept through 0,0-mode resonance, saw a peak in the transmission, saw an accompanying error signal similar to the signal shown in 5202.

3) Tried to lock. Swept the gain on the LB1005 and could not find a gain that would make it lock. Tried changing the PI-corner freq. from 10 kHz to 30 kHz and back and still could not lock.

4) Noticed that the open loop error signal displayed on the scope was DC-offset from zero. Changed the offset to zero the error signal.

5) Tried to lock again and succeeded.

6) Noticed that upon closing the loop, the error signal became offset from zero again. Turning on the integrator on the LB1005 increased DC-offset.

7) Reduced the gain on the SR560 being used as a low pass filter from 5 to 1. Readjusted the open loop error signal offset on the LB1005.

8) Closed the loop and locked. Closing the loop then caused a much smaller DC change in the signal than I had seen with the larger gain (now around 3mV). RMS fluctuations in error signal are now 1 mV (well within the linear region of the error signal).

9) Noticed transmission has a strange distorted harmonic oscillation in it a 1MHz. (Modulation freq is 230kHz, so it's not that). Checked reflected signal and also saw a strange oscillation--in a sawtooth-like pattern.

 

I intend to

1) Post oscilloscope traces here showing transmitted and reflected signal when locked.

2) Look upstream to see if the sawtooth-like oscillation is in the laser beam before it enters the cavity:

  • Sweep the temperature of the laser so that the beam is no longer resonating in the cavity.
  • Compare the reflected signal off the cavity to the signal detected before being directed into the cavity (using the PDA255 that I used for measuring the AM response of the PZT) both with and and without the frequency modulation.

3) At some point, try to close the slow digital loop, perhaps readjusting the gain.

4) Try to measure a temperature step response.

I decided to go forward and try to close the digital loop in spite of the unexplained oscillations in the transmission.

1) Plugged the 20dB attenuator into the slow port on the laser drive. This pushed the laser out of lock and, for some reason, made the laser temperature stop responding to sweeping the set point manually with the knob.

2) Plugged the output from the digital system into the slow port (with the attenuator still in place).

3) Displayed the beam seen by the camera on a monitor in the control room

4) Stepped the laser temperature using MEDM until finding the 0,1 mode. Locked to that mode.

5) Closed the digital loop (input to slow laser drive attenuated 20dB attenuator). Gain 0.010

6) Loop appeared stable for 30 minutes, then temperature began shooting off. I opened the loop, cleared history, reduced the gain to 0.008, and started it again. Loop appears stable after 15 minutes of watching. I'm going to leave it for a few hours, then come back to check on it and, if it's stable, step the can temperature.

ELOG V3.1.3-